A closed rigid container has a volume of 1 m3 and holds air at 344.8 kPa and 273 K. Heat is added until the temperature is 600 K. Determine the change of internal energy


1
Expert's answer
2021-10-05T15:43:54-0400

"V_1= 1 \;m^3 \\

p_1 = 344.8 \;kpa \\

T_1 = 273 \;K \\

T_2 = 600 \;K \\

p_1V_1 =mRT_1 \\

R= 0.287 \;kJ/kg\;K \\

344.8 \times 1 = m \times 0.287 \times 273 \\

m = 4.4007 \;kg"

The specific heat at constant volume:

"c_v = 0.718 \;kJ/kg \;K"

Internal energy \:

"ΔU = mc_vΔT \\

= 4.4007 \times 0.718 \times (T_2-T_1) \\

= 3.1597 \times (600-273) \\

= 3.1597 \times 327 \\

= 1033.2 \;kJ"

Answer: 1033 kJ